section 2: Short Answer (15 ponts) In responding to the questions in this section of the exam, please note * Please answer ALL short answer questions in this section -Be precise and succinct in your response, and Where possible you may use bullet For all these questions, to EABN FULL credit Each question in this section is worth THREE (3) points . points to provide your response. you MUST show your work Use the information below to respond to the next t Jimaine Grainger, works for GE a P Microwave Ovens, specificall hree questions that foilows Model PEB9159SFSS. The retail price of this model is $600. Best Buy, 25%. According to a an annual salary of syppiances as a Product Manager for Home and Kitchen Appliances 00,000 He has primary responsibility for the GE Wne of Built-in Inc. sells the model recent report released by the Gartner Group, the tot estimated to be 100,000 units. Jimaine estimates The report also estimates that there are l US market for bullt-in microwave ovens is that Model PEB9 1595 FSS has 12% of this market. to consumers in the US at a retail margin on the product is 400,000 potential customers in the market. Variable costs of $500,000 uring Model PEB9159SFSS are $120 while the company has invested ommission of 10% on all sales. GE Appliances typically employs the services of Regis Transport to ship all the microwave ovens to Best Buy. Regis Transport charges a flat rate of $35 per microwave that is shipped. GE Appliances has allocated $1,000,000 toward advertising and promotional expenses for this microwave oven manufact in machinery to manufacture this model. The GE Appliances salesforce is typically paid a 1) Calculate the unit contribution for Model PEB9159SFSS

1.) calculate unit contribution

2.) calculate breakeven

3.) what is the profit impact for model peb9159sfss
Answer #1

Assumption:

1) 12 % of current market share ( Not potential customers included). So number of Unit sold = 12 % x 100000 = 12000

2) Contribution , profit all re measured from GE's perspective not from best buy. So unit sales price is $600

3) Commission of sales is calculated 10% on sales price= 10% x 600 = $60

----------------------------------------------------------

Variable Cost / UNit
VC of MFg Model 120
Commision @ 10% 60
Transport Charges 35
Total Variable Cost 215

-----------------------------------------

units Sales Price Fixed Cost
12000 600 Salary 100000
Machinery 500000
Advrt 1000000
1600000

------------------------------------------------

Contribution / Unit = Sellin Price (SP) - Variable Cost / UNit(VC) = 600 - 215 = 385 / unit

--------------------------------------------

Break Even = Fixed cost / Contribution/ UNit = 1600000/385 = 4156 units

------------------------------------------------

Profit = (Selling price - Variable cost ) x no of unit sold - Fixed cost

= 385 x 12000 - 1600000

= 30,20,000